OpenStudy (anonymous):

please help!!! Simplify the expression sqrt-16/(3-3i)+(1-2i)

OpenStudy (anonymous):

a) -20-16i/9 b) 8+4i/15 c) 8-4i/15 d) -20-16i/41

OpenStudy (anonymous):

\[\frac{ \sqrt{-16} }{ 3-3i + 1 -2i }\] Combine like terms in the denominator to get: \[\frac{ \sqrt{-16} }{ -5i +4 }\] Now, since we have a negative under the radical, put an i in front and make the radical a positive 16. The square root of 16 is 4, so: \[\frac{ 4i }{ 4-5i }\]

OpenStudy (anonymous):

None of those answers look like mine. Are you sure you typed the problem accurately?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

retriceipe

OpenStudy (anonymous):

retriceipener

OpenStudy (anonymous):

azzwipe

OpenStudy (anonymous):

@danny562 retype the equation including the brackets for the sqrt. Preferably use the equation button to write it clearly. As you have it typed vinnv would be correct.

OpenStudy (anonymous):

\[\sqrt{-16} \over (3-3i) + (1-2i)\]

OpenStudy (anonymous):

if that is it then vinny is correct the only thing left to do is multiply by a factor of one (hint:complex conjugate of the denominator)

OpenStudy (anonymous):

Ah yes, I forgot about that, you would need to multiply the numerator and denominator by the conjugate of the denominator. As you can tell, my algebra is a little rusty sometimes